OpenStudy (anonymous):

Sara is reviewing recent orders at her deli to determine which meats she should order. She found that of 1,000 orders, 390 customers ordered turkey, 345 customers ordered ham, and 300 customers ordered neither turkey nor ham. Based on these data, how many of the next 1,000 customers will order both turkey and ham? Show your work and use complete sentences.

OpenStudy (anonymous):

So based on this data, we know how many people ordered turkey and how many people ordered ham correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So if the next 1000 people order the same way the first 1000 people did, how many people will order turkey?

OpenStudy (anonymous):

The answer is 735 people.

OpenStudy (anonymous):

390 will order turkey and 345 will order ham

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

@Ors I'm not sure how long you have been at Open Study but you aren't supposed to just give the answer. Help them work through it so next time they get it on their own.

OpenStudy (anonymous):

Yeah, I know.

OpenStudy (anonymous):

And @tdinatali that's correct. Now add up those two numbers and what do you get?

OpenStudy (anonymous):

735

OpenStudy (anonymous):

And there is your answer! Great job :)

OpenStudy (anonymous):

Do you understand how that works?

OpenStudy (anonymous):

that's all you have to do? i thought that there was a different probability formula that you had to find the ones that are in the intersection set instead of the individual sets. because they can order ham or turkey but how do you know if they're ordering both?

OpenStudy (anonymous):

Ohh no no you were just making it too difficult for yourself. It's just asking how many people will order either ham or turkey :)

OpenStudy (anonymous):

Math is pretty simple, but people expect it to be hard so when they come across a very easy question, they complicate it because they think if it's easy it is wrong.

OpenStudy (anonymous):

I do that time to time as well haha

OpenStudy (anonymous):

haha oh okay, thank you so much!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

@dan815 @ganeshie8 @HourglassMage @jim_thompson5910

ganeshie8 (ganeshie8):

use below : \(\large n(A \cup B) = n(A) + n(B) - n(A \cap B)\)

ganeshie8 (ganeshie8):

since 300 customers ordered neither turkey nor ham, \(n (A \cup B) = 1000 - 300 = 700\) and, you know that \(n(A) = 390\) and \(n(B) = 345\)

ganeshie8 (ganeshie8):

plug these values above and solve \(n(A \cap B)\)

ganeshie8 (ganeshie8):

\(\large 700 = 390 + 345 - n(A \cap B)\) \(\large n(A \cap B) = 390 + 345 - 700\) simplify

ganeshie8 (ganeshie8):

Yep !

OpenStudy (anonymous):

okay and now how do you find how many people will order ham and turkey? is it just 1000-35? so is it 965 people will order ham AND turkey? @ganeshie8

ganeshie8 (ganeshie8):

35 people will order ham AND turkey

ganeshie8 (ganeshie8):

\(\cap \) = AND

ganeshie8 (ganeshie8):

\(\cup \) = OR

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

I think I understand

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

35 people WILL order HAM AND TURKEY